# FlightKooba: A Fast Interpretable FTP Model

### v1

@@ Abstract
The Koopman theory is a powerful and effective modeling tool for converting nonlinear systems into linear representations, and flight trajectory prediction (FTP) is a complex nonlinear system. However, current models applying the Koopman theory to FTP tasks are not very effective, model interpretability is indeed an issue, and the Koopman operators are computationally intensive, resulting in long training times. To address this issue, this paper proposes a new modeling and control framework based on the HIPPO method, the Koopman theory, and state space equations from cybernetics: FlightKooba. Inspired by the idea of structural state space equations, FlightKooba directly constructs the Koopman operators from data. This makes the framework highly interpretable and significantly reduces the number of trainable parameters in the module, thereby greatly reducing training time. Experiments have demonstrated the superiority of the FlightKooba modeling method in terms of time and memory consumption (training time comparable to the Mamba module without using CUDA-level acceleration; memory reduced by more than 50% on most datasets, with a tenfold reduction in the number of parameters), essentially completing the FTP task. It provides a new method for the fast computation of the Koopman operators, opening up new possibilities for the combination of time series forecasting and control.
@@
Flight trajectory prediction , Koopman , Mamba , SSM , Control
journal: Neural Networks
@@
The model demonstrates excellent performance on multiple public datasets
@@ 1 Introduction
Flight trajectory prediction (FTP) is one of the fundamental research tasks in the field of aviation [6]. It is a key technology for various applications in the fields of flight operation quality assurance (FOQA) [4] and air traffic control (ATC) [33]. Flight trajectory prediction has always been an important aspect of air traffic management and safety. In flight training, trajectory prediction can be used to identify potential risks (such as attitude loss of control and altitude deviation) in advance by analyzing the trainee’s operating habits and the aircraft’s status (such as rudder deflection angle and roll rate). Flight trajectory prediction is not only a technical tool in training, but also a key factor in reshaping pilot training models. By integrating deep learning methods with physical models and combining real-time data-driven optimization strategies, we achieve comprehensive capability enhancements across the entire chain, from basic combat operations to tactical decision-making, driving the development of aviation training toward intelligence and precision. Currently, many scholars have introduced deep learning and neural networks to study flight trajectory prediction.
Deep learning methods have achieved relatively good success in modeling large complex systems using the [30] dataset. Existing deep learning technologies can be applied to various fields to predict time series data and achieve good results, such as weather forecasting [29] and traffic prediction [19].
@@
The Koopman theory [17] has attracted increasing attention in recent years and has become a promising data-driven method. By representing nonlinear systems using linear state-space models (SSM) [16], Koopman operators can convert complex nonlinear dynamic systems into linear forms. Based on this method, Williams proposed an extended dynamic modal decomposition method (EDMD) to calculate the Koopman operator [31] With the emergence and development of deep learning, Koopman theory has been combined with neural networks in many fields and has achieved great success in modeling many complex systems, such as probabilistic neural networks [14] and graph neural networks [20]. Although the prediction results of the above model are relatively accurate, it performs poorly in long sequence prediction tasks. Furthermore, neural network models themselves have a large number of parameters, and the current methods for calculating Koopman operators are time-consuming, which leads to enormous time consumption when combining the two methods.
To solve this problem, we turned our attention to the Mamba module, which is faster than the transformer module. In the Mamba structure, the core structure is similar to the linear SSM in Koopman theory. On this basis, features of recurrent neural networks (RNN) and convolutional neural networks (CNN) are integrated [7], which may also be the key to the success of the Mamba model. The core original form of this structure originates from polynomial approximation theory, namely HIPPO [8], which provides a certain theoretical basis for our discovery.
This paper proposes a time prediction model based on the HIPPO method, Koopman theory, and state space equations in cybernetics: FlightKooba. The HIPPO method was originally proposed to solve the problem of information compression in long sequences; Koopman theory solved the problem of analyzing and predicting nonlinear systems; and state space equations are mathematical tools for describing the behavior of dynamic systems. FlightKooba combines the three into one, which has great advantages in modeling complex nonlinear systems and is expected to solve the flight trajectory prediction (FTP) task. In addition, this method has a high level of interpretability that many network models lack, and it significantly reduces the number of parameters in the module, thereby greatly reducing the model training time: without a CUDA-level acceleration method, it achieves the same training time as the Mamba module. The contributions of this work are summarized as follows:
@@
    We first proposed a data-driven unsupervised modeling method called FlightKooba, which uses state space equations to link the HIPPO method and Koopman theory for modeling complex nonlinear systems.
@@
    We propose a new method for computing Koopman operators that significantly reduces the required computation time and is interpretable.
@@
    We apply the model to a flight trajectory dataset (CAFUC) to test its performance in flight trajectory prediction (FTP) tasks and obtain effective results.
@@
    We conducted experiments on other different time series datasets to validate the effectiveness and superiority of the model.
The remainder of this paper is organized as follows: Section 2 reviews the related work. Section 3 introduces FlightKooba for the FTP task. Section 4 presents the experiments, results, and visualizations. Finally, Section 5 concludes with a discussion of the findings and future work.
@@ 3 Methodology
Figure 1: Structure of FlightKooba. FlightKooba consists of two parts: the first part calculates the coefficients corresponding to the polynomials that fit the system observation function; the second part generates the corresponding Koopman operators based on the results of the first part, which are used for subsequent prediction tasks.
@@ 3.1 The Hippo Theory
In the paper “HiPPO: Recurrent Memory with Optimal Polynomial Projections,” the authors proposed the HiPPO theory, whose main idea is to use various polynomials to infinitely approximate a function in order to compress continuous information. After a series of deductions, the HIPPO algorithm ultimately arrives at an iterative formula for calculating the coefficient $c_{n}$ in front of each polynomial. In this paper, we draw on this theoretical framework and select a Legendre polynomial with good performance as the first part of FlightKooba.
The value of $c_{n}$ is related to the number of iterations, that is, it is related to time $t$. Let $(c_{0},c_{1},…,c_{n})$ be denoted as $c_{t_{k}}$, and the final result is:
@@
In the equation, $t_{k}$ denotes the $k$th $\delta_{t}$ time point. $\gamma_{k}$ denotes the function value $\gamma_{t_{k}}$ at time $t_{k}$. The matrices $\bar{N}$ and $\bar{M}$ are defined as follows:
@@
Among them, the expressions of matrices $N$ and $M$ are related to the derivation of the HIPPO algorithm. In ”HiPPO: Recurrent Memory with Optimal Polynomial Projections” the expressions of matrices $N$ and $M$ corresponding to the two methods $LegT$ and $LegS$ are as follows:
@@
In $LegT$, $\omega$ denotes the length of the historical window of interest and is a hyperparameter. The difference between $LegT$ and $LegS$ is that $LegT$ focuses more on information within a historical window length, while $LegS$ focuses more on information across the entire time series length.
@@
$c_{t_{2k+1}}=\overline{N}^{k-1}c_{t_{k}}+R\overline{M}\gamma(t_{2k})$ (8)
In the equation, $R=\left(\overline{N}^{k-1},\overline{N}^{k-2},\ldots,\overline{N},I\right)$. $\gamma(t_{2k})=(\gamma_{k+1},\gamma_{k+2},\ldots,\gamma_{2k})^{H}$ represents the function value of the observed function starting from time $2k$ and moving backward by $k-1$ steps (a total of $k$ values), i.e., a total of $k$ iterations.
@@ 3.2 The Koopman Operator
The Koopman operator is an infinite-dimensional linear operator that can transform complex nonlinear dynamical systems into linear forms. It provides data-driven analytical tools for complex systems that lack explicit mathematical models, such as fluid dynamics, molecular dynamics, and complex networks. Its core idea is to characterize changes in the state of a system through the evolution of an observation function:
@@ 3.4 The Construction Method Of FlightKooba
The coefficients $(b_{0},b_{1},…,b_{m})$ can be obtained through neural network training. At this point, the $B$ matrix will correspondingly change to:
@@
However, there is a problem with equation 29: its state variables have $n$ terms, while the right side of formula 25 has $n+1$ terms. To solve this problem, we can retain the $x_{1}$ before iteration and append it to the state vector after iteration to obtain the complete $n+1$ state variables.
@@
Figure 2: FlightKooba calculation details. The observation function is treated as a state equation and represented by Legendre polynomials. The coefficients of each polynomial are calculated using the HIPPO method, and these coefficients are then used to generate the Koopman operator, which can predict a series of future state changes.
@@
5:The matrix $\mathcal{K}$ is generated by splicing the generated $A$ and $B$ matrices;
@@ 4.1 Flight Trajectory Dataset: CAFUC
The data set used in this work, named the Multidimensional Flight Trajectory Dataset of the Civil Aviation Flight University of China (CAFUC), from the real world, was generated from realistic fixed-wing aircraft flight training at CAFUC and contains 64 parameters, four artificial labels, 150,000 total frames, 41.6 h of total flight duration in CSV format, 14,356 basic maneuvers and 168 flight training subjects. However, it contains some features with text values or no change; after removing these features, 32 features are obtained that can be used for training. The CAFUC data set is available in the following repository: https://github.com/CAFUC-JJJ/FlightKoopman.
@@ 5 Experiments
In this section, we evaluate the performance of the FlightKooba module in predicting time series. Specifically, we will evaluate (a) the convergence speed of the algorithm on different datasets with randomly initialized parameters, (b) the model accuracy of FlightKooba compared to other methods on different datasets, and (c) the training time and memory usage on different datasets with the same hyperparameter settings.
In terms of modeling performance evaluation, we compared this method with two competing methods. (1) The latest results of the original xLSTM model by the LSTM team [3] (2) The Mamba module based on the state space model (SSM) [7]. It should be noted that the FlightKooba module is designed for a module in a neural network, so we will use the mamba module and the xLSTM module for comparison.
@@ 5.1 Overall Results
For each model: We set the number of modules for each model to 2, meaning that each model consists of two stacked blocks. The batch_size is set to 32, and the seq_len is set to 8 (owing to the large amount of electricity and transportation data, the seq_len is set to 1 in both datasets to preserve the experimental results). We use the MSE loss function,the learning rate is set to 0.001, the number of epochs is set to 50, and we will evaluate the performance of each model by taking the average of 10 training runs.
For the data set: We process each data set based on the shape of the time series dataset (batch_size, features, seq_len). The first 70% of the entire data set is used as the training set. Except for the CAFUC and Lorenz datasets, which use the remaining part as the test set, the other datasets are difficult to interpret due to large fluctuations in values, so only the last 200 data points of the training set are used as the test set. (All models mentioned in this article were tested on a Linux system equipped with an RTX 4080 SUPER graphics card.)
In terms of convergence speed, FlightKooba’s convergence speed is slightly insufficient, while Mamba and xLSTM both have good convergence speeds. In terms of prediction accuracy, the three models have different advantages in different datasets, and Mamba performs best overall. In terms of training time and memory usage, FlightKooba can match the training speed of the Mamba module without CUDA-level acceleration and is even faster on some datasets.The experimental results are summarized in Figure 3.
Figure 3: MSE-time-memory bubble plots for the three models across 11 datasets (a-c). Since the text in the figure is difficult to read, the table below provides all the detailed data in the figure. As shown in Table 1.
@@
Table 1: Overview of experimental results (training set).
Datasets FlightKooba Mamba xLSTM
MSE Times Memory MSE Times Memory MSE Times Memory
CAFUC $0.0249$   0088.54 1.00 $0.0157$ $88.70$ $16.00$ 0.0094 $304.88$ $2.00$
Lorenz 0.0003 $84.44$ 1.00 $0.0005$   0084.15 $16.00$ $0.0017$ $287.31$ $2.00$
Electricity $0.0098$ 608.09 $7.00$ $0.0026$ $940.34$ $12.00$ 0.0017 $2077.96$ 6.00
ETTh1 $0.0043$ $382.51$ 1.00 0.0028 366.25 $16.00$ $0.0073$ $1331.12$ $6.00$
ETTh2 $0.0018$ $377.40$ 1.00 0.0008 344.24 $17.00$ $0.0026$ $1314.60$ $2.00$
ETTm1 $0.0010$ $1520.36$ $2.00$ 0.0007 1383.79 $17.00$ $0.0030$ $5285.54$ $4.00$
ETTm2 $0.0015$ $1506.35$ $2.00$ 0.0003 1384.00 $17.00$ $0.0010$ $5250.12$ $4.00$
Exchange $0.0081$ 166.48 1.00 0.0008 $157.06$ $2.00$ $0.0377$ $577.88$ $3.00$
ILI $0.0602$ $21.22$ 1.00 0.0050   0018.44 1.00 $0.0398$ $69.29$ $3.00$
Traffic $0.0060$ 545.94 $9.00$ 0.0022 $1398.47$ $16.00$ $0.0027$ $1878.39$ $10.00$
Weather 0.0012 $1153.25$ $3.00$ $0.0013$ 1070.23 $6.00$ $0.0014$ $4065.36$ $7.00$
The training time in the table is in $ms$ and the memory is in $GB$. The experimental results show that FlightKooba currently performs well on datasets related to physical systems, but poorly on general time series datasets. In fact, this is one of the current shortcomings of FlightKooba. However, what is even more obvious is that FlightKooba has significant advantages in terms of time and memory.
In terms of model parameters, FlightKooba has the fewest parameters. The table 2 below shows the number of parameters for all models on different datasets.
@@
datasets FlightKooba Mamba xLSTM
CAFUC 12 / 1 120 12 992 22 434
Lorenz 3 / 38 936 3 254
Electricity 62 / 33 152 711 336 1 954 775
ETTh1 5 / 84 1 764 4 709
ETTh2 5 / 84 1 764 4 709
ETTm1 5 / 84 1 764 4 709
ETTm2 5 / 84 1 764 4 709
Exchange 5 / 126 2 064 5 226
ILI 4 / 120 1 764 4 709
Traffic 52 / 87 480 4 822 028 13 444 712
Weather 14 / 420 7 140 13 403
@@
5.2 Convergence Rate
The experimental results for this section are shown in Figures 4 and 5. The Mamba module showed excellent convergence speed on different datasets, with very low loss values in the initial few epochs. xLSTM and FlightKooba performed similarly, with each having advantages on different datasets. It is worth noting that under the above parameter settings, the xLSTM model occasionally encountered a situation where the loss value became $nan$ during training, which occurred more frequently after the learning rate changed to 0.003 (possibly due to a phenomenon known as gradient explosion).
Figure 4: Convergence rate of each model on different training sets (a).
Figure 5: Convergence rate of each model on different training sets (b).
Overall, all three models converge quickly (loss values change dramatically only in the first 10 epochs on most datasets).

### v2

@@ Abstract
Flight trajectory prediction (FTP) and similar time series tasks typically require capturing smooth latent dynamics hidden within noisy signals. However, existing deep learning models face significant challenges of high computational cost and insufficient interpretability due to their complex black-box nature. This paper introduces FlightKooba, a novel modeling approach designed to extract such underlying dynamics analytically. Our framework uniquely integrates HiPPO theory, Koopman operator theory, and control theory. By leveraging Legendre polynomial bases, it constructs Koopman operators analytically, thereby avoiding large-scale parameter training. The method’s core strengths lie in its exceptional computational efficiency and inherent interpretability. Experiments on multiple public datasets validate our design philosophy: for signals exhibiting strong periodicity or clear physical laws (e.g., in aviation, meteorology, and traffic flow), FlightKooba delivers competitive prediction accuracy while reducing trainable parameters by several orders of magnitude and achieving the fastest training speed. Furthermore, we analyze the model’s theoretical boundaries, clarifying its inherent low-pass filtering characteristics that render it unsuitable for sequences dominated by high-frequency noise. In summary, FlightKooba offers a powerful, efficient, and interpretable new alternative for time series analysis, particularly in resource-constrained environments.
@@
Flight trajectory prediction , Deep learning , Koopman operator , SSM , Control
@@
The model demonstrates competitive performance on datasets exhibiting clear periodicity or physical regularities.
@@ 1 Introduction
Flight trajectory prediction (FTP) is one of the fundamental research tasks in the field of aviation [6]. It is a key technology for various applications in the fields of flight operation quality assurance (FOQA) [4] and air traffic control (ATC) [35]. Flight trajectory prediction has always been an important aspect of air traffic management and safety. In flight training, trajectory prediction can be used to identify potential risks (such as attitude loss of control and altitude deviation) in advance by analyzing the trainee’s operating habits and the aircraft’s status (such as rudder deflection angle and roll rate). Flight trajectory prediction is not only a technical tool in training, but also a key factor in reshaping pilot training models. By integrating deep learning methods with physical models and combining real-time data-driven optimization strategies, we achieve comprehensive capability enhancements across the entire chain, driving the development of aviation training toward intelligence and precision. Currently, many scholars have introduced deep learning and neural networks to study flight trajectory prediction. However, existing deep learning-based FTP models still face challenges related to computational cost and memory overhead, especially when dealing with long-term temporal dependencies.
@@
Deep learning methods have achieved relatively good success in modeling large complex systems using the [32] dataset. Existing deep learning techniques can be applied across various domains to forecast time series data with excellent results, such as weather forecasting [31] and traffic prediction [20]. Koopman theory [18] has garnered increasing attention in recent years, emerging as a promising data-driven approach. By representing nonlinear systems using linear state-space models (SSM) [17], the Koopman operator transforms complex nonlinear dynamic systems into linear forms. Building upon this approach, Williams proposed an Extended Dynamic Modal Decomposition (EDMD) method to compute the Koopman operator [33]. With the emergence and advancement of deep learning, Koopman theory has been combined with neural networks across numerous domains, achieving significant success in modeling many complex systems, such as probabilistic neural networks [15] and graph neural networks [21]. li2019.
Existing temporal data processing models (e.g., xLSTM, Mamba, etc.) or various models combined with Koopman algorithms have made tremendous progress in their respective domains. However, they all suffer from one or more of the following issues: Computational Intensity: High parameter count and slow training; Black-box nature: Difficult to interpret and trust, posing a critical flaw in safety-critical domains like aviation; Overfitting noise: These powerful models tend to learn every minor fluctuation in the data, encompassing both genuine dynamics and substantial meaningless sensor noise. In many physical systems, we are more concerned with the intrinsic core dynamics of the system than with the noise itself.
To address these issues, we propose a fundamentally different modeling approach. We contend that an excellent dynamical model should not aim to blindly fit every data point, but rather capture and predict the core patterns of system evolution. Inspired by this, we developed FlightKooba, a lightweight framework grounded in classical dynamical systems theory. It maps time series onto a set of Legendre polynomial coefficients via the HIPPO theory, and linearizes complex nonlinear dynamics using Koopman operator theory. Crucially, the system’s state transition matrix (i.e., the Koopman operator) is analytically derived from these coefficients, rather than being obtained through brute-force fitting via backpropagation. This design yields two major advantages: first, exceptionally high computational efficiency; second, clear physical and mathematical interpretability.
The main contributions of this work are as follows:
@@
    A Novel Modeling Approach: We introduce FlightKooba, a novel method for time series forecasting that shifts focus from black-box fitting to the analytical construction of dynamic operators. By integrating HiPPO, Koopman, and control theory, we provide a theoretically grounded approach for modeling underlying system dynamics.
@@
    Significant Gains in Computational Efficiency: We demonstrate that FlightKooba achieves a reduction in parameter count by several orders of magnitude compared to baseline models like Mamba and xLSTM. This directly translates to substantially faster training times and minimal memory consumption, making it ideal for resource-constrained applications.
@@
    Competitiveness on target datasets: Experiments demonstrate that for time series with strong periodic patterns or clear physical foundations, FlightKooba achieves an outstanding trade-off between performance and efficiency. It delivers competitive prediction accuracy while substantially reducing computational costs.
@@
    In-depth analysis of applicability boundaries: We conduct a rigorous analysis of the model’s theoretical limitations. We identify and explain its inherent low-pass filtering characteristics, providing clear guidance on its ideal application scenarios and elucidating why it is unsuitable for high-frequency noise-driven signals.
The remainder of this paper is organized as follows: Section 2 reviews related work. Section 3 introduces the FlightKooba framework in detail. Section 4 presents the experimental setup, results, and a dedicated discussion on the model’s limitations. Finally, Section 5 concludes the paper and outlines future research directions.
@@ 3 Methodology
Figure 1: The structure of FlightKooba. FlightKooba consists of two components: the first component (HIPPO) computes the coefficients corresponding to the polynomial that fits the system’s observation function; the second component (Kooba) generates the corresponding Koopman operator based on the results from the first component, which is used for subsequent prediction tasks. The specific matrices $M$, NN, AA, $B$, etc., shown in the figure will be detailed in the subsequent sections.
@@ 3.1 The Hippo Theory
Albert Gu proposed the HiPPO theory [8], whose main idea is to use various polynomials to infinitely approximate a function in order to compress continuous information. After a series of derivations, the HIPPO algorithm ultimately obtains an iterative formula to calculate the coefficient $c_{n}$ in front of each polynomial. In this paper, we draw on this theoretical framework and select a Legendre polynomial with good performance as the first part of FlightKooba.
The value of $c_{n}$ is related to the number of iterations, that is, it is related to time $t$ and changes with time. Let $(c_{0}(t_{k}),c_{1}(t_{k}),\cdots,c_{n}(t_{k}))$ be denoted as $c_{t_{k}}$ (the value of coefficient $c_{t_{k}}=(c_{0}(t_{k}),c_{1}(t_{k}),\cdots,c_{n}(t_{k}))$ at time $t_{k}$), and the final result is:
@@
In the equation, $t_{k}$ denotes the $k$th $\Delta t$ time point, where $\Delta t$ is a hyperparameter (for convenience of calculation, if the sequence length is $T$, we set $\Delta t$ to $\frac{1}{T}$). $\gamma_{k}$ denotes the function value $\gamma(t_{k})$ at time $t_{k}$, i.e., the actual recorded data sample point. The $n$ in $c_{n}$ denotes the number of terms used in the polynomial fitting of the objective function. It is important to note that $c_{n}$ represents the coefficient corresponding to the $n$th term when fitting the function, while $c_{t_{k}}$ denotes the corresponding value of the coefficient vector $c_{t_{k}}=(c_{0}(t_{k}),c_{1}(t_{k}),\cdots,c_{n}(t_{k}))$ at time $t_{k}$. Finally, the matrices $\overline{N}$ and $\overline{M}$ are defined as follows:
@@
Among them, the expressions of the matrices $N$ and $M$ are related to the derivation of the HIPPO algorithm. In the original text [8], the expressions of matrices $N$ and $M$ corresponding to the two methods $LegT$ and $LegS$ are as follows:
@@
In $LegT$, $\omega$ denotes the length of the historical window of interest (in subsequent experiments, we will assign this value to seq_len), which is a hyperparameter. The difference between $LegT$ and $LegS$ is that $LegT$ focuses more on information within a historical window length, while $LegS$ focuses more on information across the entire time series length.
@@
$c_{t_{2k+1}}=\overline{N}^{k-1}c_{t_{k}}+R\overline{M}\gamma_{t_{2k}}$ (8)
In the equation, $R=\left(\overline{N}^{k-1},\overline{N}^{k-2},\cdots,\overline{N},I\right)$. $\gamma_{t_{2k}}=(\gamma_{k+1},\gamma_{k+2},\cdots,\gamma_{2k})^{T}$ denotes the function value of the observation function starting from time $2k$ and moving backward by k1k-1 steps (a vector with $k$ values, which allows both $c_{t_{k}}$ and $\gamma_{t_{2k}}$ to represent vectors for clarity), that is, a total of $k$ steps are iterated. Detailed derivation of this multistep iteration method is provided in Appendix B.
@@ 3.2 The Koopman Operator
The Koopman operator is an infinite-dimensional linear operator that can transform complex nonlinear dynamical systems into linear forms. Provides data-driven analytical tools for complex systems that lack explicit mathematical models, such as fluid dynamics, molecular dynamics, and complex networks. Its core idea is to characterize changes in the state of a system through the evolution of an observation function:
@@ 3.4 The Construction Method Of FlightKooba
The coefficients $(b_{0},b_{1},\cdots,b_{m})$ can be obtained through neural network training. At this point, the $B$ matrix will correspondingly change to:
@@
After obtaining the analytical forms of matrices $A$ and $B$, we can utilize the state transition equation in SSM to compute the state after $k$ steps in actual calculations. Details regarding this process can be found in the appendix. However, there is a problem with equation 29: its state variables have $n$ terms, while the right side of formula 25 has $n+1$ terms. To solve this problem, we can retain the $x_{1}$ before iteration and append it to the state vector after iteration to obtain the complete $n+1$ state variables.
@@
Figure 2: Within the framework of flight trajectory prediction applications, FlightKooba computes details. The observation function serves as the state equation, represented by Legendre polynomials. The HIPPO method calculates the coefficients for each polynomial, which are then used to generate the Koopman operator. This operator can predict a sequence of future state changes.
@@
5:The matrix $\mathcal{K}$ is generated by splicing the generated AA and $B$ matrices;
@@ 4.1 Flight Trajectory Dataset: CAFUC
The data set used in this work, named the Multidimensional Flight Trajectory Dataset of the Civil Aviation Flight University of China (CAFUC), from the real world, was generated from realistic fixed-wing aircraft flight training at CAFUC and contains 64 parameters, four artificial labels, 150,000 total frames, 41.6 h of total flight duration in CSV format, 14,356 basic maneuvers and 168 flight training subjects. However, it contains some features with text values or no change; after removing these features, 12 features are obtained that can be used for training. The CAFUC data set is available in the following repository: https://github.com/CAFUC-JJJ/FlightKoopman.
@@ 5 Experiments
In this section, we evaluate the performance of the FlightKooba module in forecasting time series. Specifically, we will assess the following under identical hyperparameter settings: (a) the training time and memory consumption of different models across various datasets, and (b) the accuracy demonstrated by different models on these datasets.
For modeling performance evaluation, we compare the modules obtained by our method with two additional modules; furthermore, we compare with two full modeling approaches. Modules: (1) The latest xLSTM model from the LSTM team [3], (2) The Mamba module based on the State Space Model (SSM) [7]; Full Methods: We used two state-of-the-art modeling approaches for time series or flight trajectory prediction (Koopa and FlightBERT++) [23, 10].
It should be noted that xLSTM is categorized as a module because the official code interface allows users to select the number of blocks within xLSTM, whereas the latter two methods directly provide complete network designs. However, the FlightKoopa module is designed as a single module within a neural network. Therefore, comparisons with the Mamba module and xLSTM module are the primary focus, while experimental results comparing it with the subsequent two complete methods serve only as reference.
@@ 5.1 Overall Results
For each model: We set the number of modules for each model to 2, which means that each model consists of two stacked blocks. The batch_size is set to 32 and the seq_len is set to 8. We use the MSE loss function,the learning rate is set to 0.001, the number of epochs is set to 50, and we will evaluate the performance of each model by taking the average of 10 training runs.
For the data set: We process each data set based on the shape of the time series data (batch_size, seq_len, features). The first 70% of the entire data set is used as a training set, followed by the next 200 data points as a test set. (All models mentioned in this paper were tested on Linux systems equipped with RTX 5880 Ada graphics cards.)
In terms of prediction accuracy: Koopa and FlightBERT++ demonstrated the best performance, achieving exceptionally low loss values on the training set. This also highlights the advantages of full neural network modeling approaches in various tasks. However, the training time and memory overhead associated with such modeling methods cannot be overlooked. Regarding training time and memory usage: The FlightKooba module exhibited the shortest training time and the lowest memory consumption among the five models. The experimental results (including MSE, training time, and memory consumption) are shown in Figure 3(a), 3(b) and 3(c).
(a) MSE-time-memory bubble plots for FlightKooba across 11 datasets.
(b) MSE-time-memory bubble plots for Mamba and xLSTM across 11 datasets.
(c) MSE-time-memory bubble plots for Koopa and FlightBERT++ across 11 datasets.
@@
Table 1: Module comparison experiment results (training set).
Datasets FlightKooba Mamba xLSTM Koopa FlightBERT++
MSE Times Memory MSE Times Memory MSE Times Memory MSE Times Memory MSE Times Memory
CAFUC 0.0029 2768.27 1.00 0.0008 5004.75 1.00 0.0006 13490.10 1.00 0.0002 15755.00 1.00 0.0003 12481.30 1.00
Lorenz 0.0002 327.61 1.00 0.0005 645.60 1.00 0.0001 1156.78 1.00 5.9e-7 2109.12 1.00 0.0004 1698.95 1.00
Electricity 0.0058 2240.92 2.00 0.0026 4556.50 3.00 0.0026 7665.13 2.00 0.0007 14284.80 3.00 0.0030 11303.70 2.00
ETTh1 0.0044 1474.76 1.00 0.0018 2857.34 2.00 0.0021 5074.68 2.00 0.0006 9303.71 2.00 0.0004 7466.64 2.00
ETTh2 0.0025 1468.62 1.00 0.0009 2850.76 1.00 0.0008 5096.36 2.00 0.0002 9285.73 2.00 0.0049 7435.49 2.00
ETTm1 0.0010 4684.61 1.00 0.0007 8724.79 1.00 0.0008 15402.40 2.00 0.0002 28284.90 2.00 0.0002 22436.80 2.00
ETTm2 0.0016 4775.36 1.00 0.0004 8743.13 1.00 0.0003 15419.30 2.00 7.0e-5 28289.50 2.00 0.0026 22402.2 2.00
Exchange 0.0071 652.31 1.00 0.0029 1234.16 1.00 0.0038 2211.59 2.00 3.2e-5 4025.54 2.00 0.0075 3225.84 2.00
ILI 0.0062 81.29 1.00 0.0015 150.49 1.00 0.0036 269.47 2.00 0.0005 489.97 2.00 0.0054 392.60 1.00
Traffic 0.0042 2532.24 2.00 0.0021 3296.00 4.00 0.0024 5136.43 3.00 0.0010 9774.44 4.00 0.0023 7452.31 3.00
Weather 0.0012 3289.06 1.00 0.0011 6019.41 2.00 0.0013 10446.30 2.00 0.0002 19238.60 2.00 0.0009 15189.00 2.00
The training time in the table is in $ms$ and the memory is in $GB$. Experimental results demonstrate that FlightKooba exhibits significant advantages in terms of memory and computational overhead across all datasets. The lowest mean squared error (MSE) values are achieved by the Koopa and FlightBERT++ models. This is because these two approaches employ full-fledged models with a large number of trainable parameters, resulting in superior fitting performance. In contrast, the first three methods merely stack two modules together, and therefore, their insufficient fitting capability is expected.
Once again, it is important to reiterate that FlightKooba was designed as a module capable of being inserted into any network layer, with optimization targets focused on minimizing memory and computational overhead. This approach inevitably involves some trade-offs in accuracy. Although MSE performance in the training set is slightly inferior, this does not necessarily indicate that the FlightKooba module is ineffective. In contrast, in subsequent dimension-wise predictions, the curves fitted by the FlightKooba module did not show a significant performance gap. (section 5.2)
FlightKooba exhibits minimal memory and time overhead during training because we perform extreme compression of trainable parameters during model construction. The table below 2 displays the number of parameters for all models in different datasets.
@@
datasets FlightKooba Mamba xLSTM Koopa FlightBERT++
CAFUC 4 / 320 66,892 65,172 122,040 182,353
Lorenz 3 / 38 65,860 64,140 113,816 178,769
Electricity 62 / 40,404 106,753 105,033 439,692 320,785
ETTh1 5 / 84 66,247 64,527 116,900 180,113
ETTh2 5 / 84 66,247 64,527 116,900 180,113
ETTm1 5 / 84 66,247 64,527 116,900 180,113
ETTm2 5 / 84 66,247 64,527 116,900 180,113
Exchange 5 / 126 66,376 64,656 117,928 180,561
ILI 4 / 120 66,247 64,527 116,900 180,113
Traffic 52 / 110,160 176,542 174,822 995,840 563,153
Weather 14 / 420 68,053 66,333 131,292 186,385